OpenStudy (anonymous):

Kinda confused with this one . Show me how to do it ? : ) Find a number whose product with 6 is the same as its sum with 60. (Hint: you may want to try guess and check or a table.) Show all steps or reasoning. Be sure to check your answer for reasonableness.

OpenStudy (anonymous):

Set it up like this: 6x=x+60 (6 times the number is the same (equals) the number (x) plus 60 Then solve for x: 6x=x+60 6x-x=60 (subtract x from both sides) 5x=60 (6x - 1x = 5x on the left side) x=60/5 (divide both sides by 5 to get the x by itself) x=12 And that's it!

OpenStudy (anonymous):

Oh , it makes so much sense now : ) Thankyou <3